OpenStudy (anonymous):

I need to solve this quadratic equation but having problems finding the domain, range and how to graph: f(x)=(x+4)^2-9

OpenStudy (anonymous):

domain: values that x can be without causing the equation to be false in this case the values of x are all real numbers, so \[domain: (-\infty,\infty)\] the range is the values of y or f(x) that are evaluated from each value of x in the domain. This means that the range depends on the domain. so if the domain is infinity the range wil be as well do you have a graphing calculator? if not: x^2 is a parabola, with vertex on (0,0). (x+4)^2 is a parabola with vertex on (0,0) but the shape of the parabola is different f(x) = (x+4)^2 - 9 means the (x+4)^2 parabola has been translated up 9 in the y-direction. So this parabola should have its vertex on (0,9) For the rest of the points, plug in reasonable values of x into the f(x) equation and you should be able to graph.